Calculate the heat change associated with cooling a 400g iron bar from 126C to 40C. ( Cp= .44)
1
Expert's answer
2020-10-01T07:52:15-0400

Solution:

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature

Ti = initial temperature


Thus:

q = (400 g) × (0.44 J oC-1 g-1) × (40 - 126)oC = -15136 J = -15.14 kJ

q = -15.14 kJ

The heat change = IqI = 15.14 kJ


Answer: The heat change is 15.14 kJ.